Determine the electric field produced on any point in space by a very long line


(infinite) charged with a uniform density λ

1
Expert's answer
2022-04-04T09:04:09-0400

"\phi=\oint E.dA"


"\phi=\oint _{s_1}E.dA+\oint _{s_2}E.dA+\oint _{s_3}E.dA"

"\phi=0+0+\smallint E.dA"


"\phi=E(2\pi r)\\2\pi xrE =\frac{x\lambda}{\epsilon}"

"E=\frac{\lambda}{2\pi\epsilon r}"